Maurice is driving his two dogs to the park. One is called Gaston and has a mass of 72 Kg.the second is called Belle and has a mass of 24kg. When belle sits on the car it compresses by 0.22m. The total spring constant is 1,100 N m-1. Calculate the elastic potential energy stored in the spring.
1
Expert's answer
2018-05-10T15:47:08-0400
Elastic potential energy can be found using formulae: U=(cx^2)/2, where c is spring constant, x is compression.
We can evaluate energy: U=(cx^2)/2=(1100*〖0.22〗^2)/2=26.62 J.
Answer: 26.62 J.